OpenStudy (anonymous):

solve log(v-1/v-16)=2

OpenStudy (anonymous):

i gotta know the steps to do it on paper

OpenStudy (anonymous):

log base ten, so what we could do is simply: raise both sides to 10, thereby elminating the log: and leaving us with: (v-1/v-16)=10^2 then im sure you can isolate v byyourself

OpenStudy (anonymous):

i get 1599/99 as my answer but the correct answer is 533/33

OpenStudy (anonymous):

\[\log(v-1)=2+\log(v-16)\] \[\log(v-1)-\log(v-16)=2\] \[\log \left(\begin{matrix}v-1 \\ v-16\end{matrix}\right)=2\] \[v-1=100v-1600\]

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

okay, so whats the problem

OpenStudy (anonymous):

v=100v-1599 99v=-1599 v=1599/99

OpenStudy (anonymous):

\[99v=1599 =1599/99 so its \not 533/33\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but the answer is 533/33

OpenStudy (anonymous):